Eugenol, a compound found in nutmeg and cloves, has the formula C10H12O2. What is the boiling point of a solution containing 0.281 g of this compound dissolved in 10.0 g of benzene? The normal boiling point of benzene is 80.10 °C, and its boiling point elevation constant is 2.53 °C/m.
Answer #1

Let compound be X

Mole of X = 0.281/164 = 0.001713 mol

Molality (m) = mole of X /mass of benzene (in kg)

= 0.00171/0.01 = 0.1713 m

Now elevation in boiling point

Delta T = kb × m

T-80.1 = 0.1713× 2.53

T = 80.533 °C

Final boiling point is 80.53°C